Johnny Come Lately Computers (JCMC) has 20,000 bonds outstanding, selling at 98 percent of par; 300,000 shares of stock outstanding, selling at $32.71 per share; and 50,000 shares of preferred stock, trading at $82 per share. What weight will common stock have in the computation of the firm's WACC?
Answer #1

Total value of bonds=20000*(1000*98%)=$19.6 million

Total value of common stock=(300,000*32.71)=$9.813 million

Total value of preferred stock=(50000*82)=$4.1 million

Total value=(19.6+9.813+4.1)=$33.513 million

Hence weight of common stock=9.813/33.513

=29.28%(Approx).